# A Comprehensive Benchmarking Analysis of Fault Recovery in Stream Processing Frameworks

### v1

@@ 9.1. Key Observations and Takeaways
  • In use cases where at-least-once semantics guarantees are required, Flink provides currently the best recovery, which is both fast and stable. Spark Structured Streaming can also provide consistent fault recovery for applications where latency is not a critical requirement. Kafka Streams takes longer to recover from failures and presents volatile behavior.
@@
  • Flink is the best candidate when exactly-once semantics guarantees are necessary. On the other hand, Kafka Streams has limited performance, and Spark Structured Streaming has limited native support for EOS.
@@ 9.2. Results Implications
The fault recovery capabilities should be a criterion for choosing a stream processing framework. Therefore, one of the main implications of our findings is that Kafka Streams, despite its growing interest, needs improvements in terms of fault recovery. This implication is relevant because our findings indicate that in contrast to what was reported in the past (van Dongen and van den Poel, 2021), Kafka Streams’ unstable recovery can compromise QoS and service level agreements. The finding that Kafka Streams faces performance instabilities after failures due to its current repartitioning strategy has relevant implications. We intend to share our findings with Kafka Streams’ community and contribute with future enhancements.
@@
With the significant impact of fault recovery on applications’ performance and QoS, we argue that modern performance benchmarking should be extended to support more realistic scenarios. We believe that performance is still a major goal, but it should be continuously improved. This would be a starting point to better representing real-world scalable computing deployments in clouds.
Fault recovery has a high impact on advanced optimizations necessary in modern stream processing, such as run-time self-adaptations for auto-scaling. This is because applying the runtime modifications involves recovery. Therefore, our findings are also directly applicable to auto-scaling. We also believe that our experimental method is generic enough to be useful for evaluating other domains. This is mostly because our method is comprehensive and includes insights into how chaos engineering can be applied to long-running computing systems.
@@ 9.3. Limitations and Threats to Validity
We expect all threats to the internal validity discussed in our previous works to also apply to our results (Henning et al., 2024). In addition to known threats to external validity, the fault measurements have additional configurations and executions threats. The failure injection interval was defined according to insights from real-world industry deployments, where there can be hundreds or thousands of worker instances, thus increasing failure likelihood. However, we still need more comprehensive statistical studies to provide insights into failure rates in large-scale cloud deployments. From our standpoint, we believe that new comprehensive studies are needed to characterize the failure rates in modern public hyperscalers, similar to what we have seen in high-performance computing centers (Schroeder and Gibson, 2009).
@@ 9.4. Future Work
We believe that future analyses to extend ours are still needed. A relevant lesson learned from the experiments is that 5 minutes between two injected failures is not enough to measure the recovery time in all scenarios. Therefore, future work is to experiment with failures injected over longer intervals. We also believe the following are necessary complementary analyses:
@@ Acknowledgements.
We would like to thank Johannes Kepler University Linz and Dynatrace for co-funding this research. We thank Bernhard Kepplinger for providing insights on the method for failure detection.

### v2

@@ 9.1. Observations Summary and Takeaways
  • In use cases where at-least-once guarantees are suitable, Flink and Spark Structured Streaming provide a consistently fast and stable recovery. Spark Structured Streaming can be a suitable alternative when low latency is not required. On the other hand, Kafka Streams takes longer to recover from failures and has volatile behavior.
@@
  • Flink is the most suitable framework for use cases requiring exactly-once semantics (EOS) guarantees, providing fast recovery and stable performance. Conversely, Kafka Streams’ executions have lower performance when EOS is enabled, and Spark Structured Streaming has limited native support for guaranteeing EOS in executing applications.
9.2. Findings discussion
Although many of the insights we discussed regarding how stream processing frameworks recover from failures align with related literature (van Dongen and van den Poel, 2021; Wang et al., 2021, 2022), our experiments reveal new trends and insights that warrant consideration. To the best of our knowledge, the following are key novel findings:
  • $\star$
    Finding: Kafka Streams provides limited fault recovery.
  • Rationale: The current Kafka Streams’ partition assignment strategy has not been thoroughly evaluated in related approaches (van Dongen and van den Poel, 2021; Wang et al., 2021). Kafka Streams was only tested under failures in (van Dongen and van den Poel, 2021) with a single failure injected. Our experimental method reached this finding due to the recurrent injection of different failures and recovery time measurements.
  • $\star$
    Finding: Flink currently provides fast and stable recovery.
  • Rationale: Measuring the stability across recurrent failures is a novelty of our experimental method, which also covers more recent enhancements implemented on Flink. Relevant enhancements in Flink include the reactive mode, as well as optimizations in checkpointing and load management (Fragkoulis et al., 2024).
10. Conclusions
Ensuring fault tolerance in data-intensive, event-driven applications is crucial for successful industry deployments. In this paper, we contribute with a more comprehensive evaluation method and fault recovery analysis. Such an evaluation was enabled by ShuffleBench’s (Henning et al., 2024) flexible deployment, parameters, and extension to fault recovery metrics.
@@ 10.1. Results Implications
One of the main implications of our findings is that Kafka Streams, despite its growing interest, needs improvements in terms of fault recovery. This implication is relevant because our findings indicate that in contrast to what was reported in the past (van Dongen and van den Poel, 2021), Kafka Streams has unstable recovery that can compromise QoS and service level agreements. The finding that Kafka Streams faces performance instabilities after failures due to its current rebalancing strategy has relevant implications. We intend to share our findings with Kafka Streams’ community and contribute with future enhancements.
@@
The insights from our experimental results can also impact on advanced optimizations necessary in modern stream processing, such as run-time self-adaptations for auto-scaling. This connection arises because implementing run-time adaptations in distributed executions inherently involves fault recovery procedures. Considering that our experimental method is comprehensive and includes insights into how chaos engineering can be applied to long-running computing systems, it is expected to be generic enough to be useful for evaluating other domains.
@@ 10.2. Limitations and Threats to Validity
We expect all threats to the internal validity discussed in our previous works to also apply to our results (Henning et al., 2024). In addition to known threats to external validity, the fault measurements have additional configurations and execution threats. The failure injection interval was defined according to insights from real-world industry deployments, where there can be hundreds or thousands of worker instances, thus increasing failure likelihood. However, we still need more comprehensive statistical studies to provide insights into failure rates in large-scale cloud deployments. From our standpoint, we believe that new comprehensive studies are needed to characterize the failure rates in modern public hyperscalers, similar to what we have seen in high-performance computing centers (Schroeder and Gibson, 2009).
@@ 10.3. Future Work
We expect that future analyses to extend ours are still needed. With the significant impact of fault recovery on applications’ performance and QoS, we argue that modern performance benchmarking should be extended to support more realistic scenarios. Performance is still a major goal, but it should be continuously improved. This would be a starting point to better representing real-world scalable computing deployments in clouds.
A relevant lesson learned from the experiments is that 5 minutes between two injected failures is not enough to measure the recovery time in all scenarios. Therefore, future work is to experiment with failures injected over longer intervals. We also believe the following are necessary complementary analyses:
@@
  • Extend the measurements to cover execution costs with the stream processing frameworks.
@@ Acknowledgements.
We would like to thank Johannes Kepler University Linz and Dynatrace for co-funding this research. We thank Bernhard Kepplinger for providing insights into the method for automating the measurement of failure recovery times.